# Tail Distribution of Regret in Optimistic Reinforcement Learning

### v1

@@ 4  Main Results
It is natural to compare Corollary 1 with the best existing gap-dependent bounds for tabular MDPs. For model-based algorithms, simchowitz2019non show that, given a global $Q^{\star}$-gap, for any fixed confidence level $\delta\in(0,1)$, the cumulative regret satisfies the following high-probability bound:
$\mathbb{P}\left(R_{K}\geq\tilde{\mathcal{O}}\!\Big(H^{4}SA(S\vee A)+\frac{H^{3}SA}{\texttt{gap}^{\star}}\log(K/\delta)\Big)\right)\;\leq\;\delta.$
@@
$\mathbb{E}[R_{K}]\;\leq\;\tilde{\mathcal{O}}\!\Big(\frac{H^{3}SA}{\texttt{gap}^{\star}}\log K\Big).$
Similar instance-dependent rates have been obtained for optimistic $Q$-learning and related model-free methods (yang2020q; velegkas2022reinforcement). All the algorithms presented in these works depend on the specific choice of confidence level $\delta\in(0,1)$.
On the contrary, our results are instance-dependent in a comparable regime, but they are organized around the entire tail rather than a single quantile. Instead of working at a fixed confidence level $\delta$ and producing a bound, Corollary 1 describes how the entire curve $x\mapsto\mathbb{P}(R_{K}\geq x)$ behaves: for $x$ above the instance-dependent baseline $m_{K}$ we obtain a sub-Gaussian tail
$\mathbb{P}(R_{K}\geq x)\;\leq\;\exp\!\Big(-\tilde{\mathcal{O}}\big(x^{2}/(H^{3}K)\big)\Big),$
up to a transition scale that depends on $(H,K,\alpha)$, and beyond that scale the tail becomes sub-Weibull, with decay of order $\exp\{-\tilde{\mathcal{O}}(K^{\alpha})\}$ in the KD case or $\exp\{-\tilde{\mathcal{O}}(x^{\alpha})\}$ in the KI case. Notice that if we fix a confidence level $\delta=c\delta_{K}$ for some $c>1$, and invert our tail bound, we recover a high-probability guarantee of the form
$\mathbb{P}\!\left(R_{K}\;\geq\;\tilde{\mathcal{O}}\big(H^{4}SA(S+K^{\alpha})/\texttt{gap}^{*}+\sqrt{H^{3}K\log(1/\delta)}\big).\right)\;\leq\;\delta.$
The key difference lies in the dependence on $K$: our bound incurs a larger $K$-dependence due to the more aggressive exploration strategy required in order to obtain a tail bound that is valid simultaneously for all $x>\mathbb{E}[R_{K}]$.
@@ 5  Proof Analysis
Recall that $\bm{\pi}^{*}=\{\pi^{*}_{h}\}_{h=0}^{H-1}\in{\bf\Pi}^{*}$, where $\bf\Pi$$^{*}$ denotes the set of optimal policies for the transition kernel $\bm{P}^{*}=\{P^{*}_{h}\}_{h=0}^{H-2}$. At episode $k$, the algorithm computes a greedy policy $\pi^{k}=\{\pi^{k}_{h}\}_{h=0}^{H-1}$ using an optimistic estimate of the optimum value function $\{V^{*}_{h}\}_{h=0}^{H-1}$ based on data collected up to episode $k$. Recall that $R_{K}$ denote the cumulative regret incurred up to episode K1K-1. Then we have:
$R_{K}$ $=\sum_{k=0}^{K-1}\left(V_{0}^{\bm{\pi}^{*}}(s_{0}^{k})-V_{0}^{\bm{\pi}^{k}}(s_{0}^{k})\right)$
$=\sum_{k=0}^{K-1}\left(V_{0}^{\bm{\pi}^{*}}(s_{0}^{k})-V_{0}^{\bm{\pi}^{k}}(s_{0}^{k})\right)\mathbbm{1}\{\bm{\pi}^{k}\notin\mathbf{\Pi}^{*}\}$
$=\underbrace{\sum_{k=0}^{\lceil K^{\gamma}\rceil-1}\left(V_{0}^{\bm{\pi}^{*}}(s_{0}^{k})-V_{0}^{\bm{\pi}^{k}}(s_{0}^{k})\right)\mathbbm{1}\{\bm{\pi}^{k}\notin\mathbf{\Pi}^{*}\}}_{R_{K}^{(0)}}$
$\qquad\qquad+\underbrace{\sum_{k=\lceil K^{\gamma}\rceil}^{K-1}\left(V_{0}^{\bm{\pi}^{*}}(s_{0}^{k})-V_{0}^{k}(s_{0}^{k})\right)\mathbbm{1}\{\bm{\pi}^{k}\notin\mathbf{\Pi}^{*}\}}_{R_{K}^{(1)}}$
$\qquad\qquad+\underbrace{\sum_{k=\lceil K^{\gamma}\rceil}^{K-1}\left(V_{0}^{k}(s_{0}^{k})-V_{0}^{\bm{\pi}^{k}}(s_{0}^{k})\right)\mathbbm{1}\{\bm{\pi}^{k}\notin\mathbf{\Pi}^{*}\}}_{R_{K}^{(2)}}.$
Here, $R_{K}^{(0)}\leq\lceil K^{\gamma}\rceil H$ denotes the initial regret incurred during episodes with insufficient data. The term $R_{K}^{(1)}$ corresponds to the optimistic estimation error, that is, the discrepancy between the optimal value under the true model $\bm{P}^{*}$ and the value estimated by the optimistic algorithm. Finally, $R_{K}^{(2)}$ captures the deviation between the optimistic estimate of the value of policy $\pi^{k}$ and its true value function.
@@
We begin by analyzing the term $R_{K}^{(1)}$. The bonus function $b_{h}^{k}(\cdot)$ is chosen so that $V_{h}^{k}$ provides an optimistic estimate of $V_{h}^{*}$, provided that the estimated transition kernel $\hat{P}_{h^{\prime}}^{k}$ and the true transition kernel $P_{h^{\prime}}^{*}$ remain close for all horizons $h^{\prime}\geq h$. To formalize this, we define the “good set” of estimation events at episode $k$ as
$\mathcal{G}_{k}=\bigcap_{h=0}^{H-1}\bigcap_{s,a}\left\{\|P_{h}^{*}(\cdot|s,a)-\hat{P}_{h}^{k}(\cdot|s,a)\|_{1}\leq\frac{2b_{h}^{k}(n_{h}^{k}(s,a))}{H-h-1}\right\}.$
@@
Under the event $\mathcal{G}_{k}$, we claim that $Q_{h}^{k}(s,a)\geq Q_{h}^{*}(s,a)$ for all $(s,a,h)\in\mathcal{S}\times\mathcal{A}\times[H]$. We establish this by induction on the horizon. At the terminal horizon, we have $Q_{H-1}^{k}(s,a)=Q_{H-1}^{*}(s,a)$ for all $(s,a)\in\mathcal{S}\times\mathcal{A}$. Assuming $Q_{h+1}^{k}(s,a)\geq Q_{h+1}^{*}(s,a)$ for all $(s,a)$, we now show that the same inequality holds for horizon $h$.
@@
$\qquad\geq-b_{h}^{k}(n_{h}^{k}(s,a)),$ (By definition of $\mathcal{G}_{k}$)
where $c=\tfrac{1}{2}\big(\max_{x}V_{h+1}^{*}(x)+\min_{x}V_{h+1}^{*}(x)\big),$ $\mathrm{span}(f)=\max_{x}f(x)-\min_{x}f(x)$, and $\mathbf{1}$ denotes the all-ones vector. Hence, under $\mathcal{G}_{k}$, taking maxima and recursing yields
$Q_{h}^{k}(s,a)\geq Q_{h}^{*}(s,a),\qquad V_{h}^{k}(s)\geq V_{h}^{*}(s),$
for all $(s,a,h)\in\mathcal{S}\times\mathcal{A}\times[H]$. This implies that under the event $\cap_{k=\lceil K^{\gamma}\rceil}^{K-1}\mathcal{G}_{k}$, we have $R_{K}^{(1)}\leq 0$.
Next, we bound the probability of the bad event $\left(\cap_{k=\lceil K^{\gamma}\rceil}^{K-1}\mathcal{G}_{k}\right)^{\mathrm{c}}$:
$\mathbb{P}\left(\left(\cap_{k=\lceil K^{\gamma}\rceil}^{K-1}\mathcal{G}_{k}\right)^{\mathrm{c}}\right)$
@@
$\leq\sum_{s,a,h}\sum_{n=1}^{K}2^{S}\exp\left(-\frac{n}{2}\left(\frac{2b_{h}^{n\vee\lceil K^{\gamma}\rceil}(n)}{H-h-1}\right)^{2}\right)$ ((weissman2003inequalities, Theorem 2.1))
$\leq SAH\begin{cases}\displaystyle\ K\exp\left(-2\mu K^{\alpha}\right),&\text{{KD}},\\[8.61108pt] \displaystyle K\exp\left(-2\mu K^{\gamma\alpha}\right),&\text{{KI}}.\end{cases}$ (Lemma 1)
@@
Decomposing the tail probability of the regret into the contribution of the bad estimation event and its complement, we have
$\mathbb{P}(R_{K}\geq x)$ $\leq\mathbb{P}(R_{K}^{(1)}+R_{K}^{(2)}\geq x-\lceil K^{\gamma}\rceil H)$
@@
$\leq\mathbb{P}\left((\cap_{k=\lceil K^{\gamma}\rceil}^{K-1}\mathcal{G}_{k})^{\mathrm{c}}\right)+\mathbb{P}\left(\left\{R_{K}^{(2)}\geq x-\lceil K^{\gamma}\rceil H\right\}\cap(\cap_{k=\lceil K^{\gamma}\rceil}^{K-1}\mathcal{G}_{k})\right),$
since $R_{K}^{(1)}\leq 0$ under $\cap_{k=\lceil K^{\gamma}\rceil}^{K-1}\mathcal{G}_{k}$.
Next, we bound $R_{K}^{(2)}=\sum_{k=\lceil K^{\gamma}\rceil}^{K-1}\Delta_{k,0}^{(2)}(s_{0}^{k})$, under the event $\cap_{k=\lceil K^{\gamma}\rceil}^{K-1}\mathcal{G}_{k}$, where $\Delta_{k,0}^{(2)}(s_{0}^{k})=\left(V_{0}^{k}(s_{0}^{k})-V_{0}^{\bm{\pi}^{k}}(s_{0}^{k})\right)\mathbbm{1}\{\bm{\pi}^{k}\notin\mathbf{\Pi}^{*}\}$. We express $\Delta_{k,0}^{(2)}(s_{0}^{k})$ recursively, introducing terms that can be bounded either via concentration inequalities for $\hat{\bm{P}}_{k}$ around $\bm{P}^{*}$ or by the bonus functions $\{b_{h}^{k}(\cdot)\}_{h}$.
$\Delta_{k,0}^{(2)}(s_{0}^{k})$ $\leq b_{0}^{k}(n_{0}^{k}(s_{0}^{k},a_{0}^{k}))\mathbbm{1}\{\bm{\pi}^{k}\notin\mathbf{\Pi}^{*}\}$
$\qquad+\sum_{s^{\prime}}\left(\hat{P}^{k}_{0}(s^{\prime}\mid s_{0}^{k},a_{0}^{k})-P^{*}_{0}(s^{\prime}\mid s_{0}^{k},a_{0}^{k})\right)V_{1}^{k}(s^{\prime})\mathbbm{1}\{\bm{\pi}^{k}\notin\mathbf{\Pi}^{*}\}$
$\qquad+\sum_{s^{\prime}}P^{*}_{0}(s^{\prime}\mid s_{0}^{k},a_{0}^{k})\left(V_{1}^{k}(s^{\prime})-V_{1}^{\bm{\pi}^{k}}(s^{\prime})\right)\mathbbm{1}\{\bm{\pi}^{k}\notin\mathbf{\Pi}^{*}\}$
$\leq b_{0}^{k}(n_{0}^{k}(s_{0}^{k},a_{0}^{k}))\mathbbm{1}\{\bm{\pi}^{k}\notin\mathbf{\Pi}^{*}\}$
$\qquad+\frac{H-1}{2}\|P_{0}^{*}(\cdot\mid s_{0}^{k},a_{0}^{k})-\hat{P}_{0}^{k}(\cdot\mid s_{0}^{k},a_{0}^{k})\|_{1}\mathbbm{1}\{\bm{\pi}^{k}\notin\mathbf{\Pi}^{*}\}$
$\qquad+\sum_{s^{\prime}}P^{*}_{0}(s^{\prime}\mid s_{0}^{k},a_{0}^{k})\Delta_{k,1}^{(2)}(s^{\prime})$
$\leq 2b^{k}_{0}(n_{0}^{k}(s_{0}^{k},a_{0}^{k}))\mathbbm{1}\{\bm{\pi}^{k}\notin\mathbf{\Pi}^{*}\}+\sum_{s^{\prime}}P^{*}_{0}(s^{\prime}\mid s_{0}^{k},a_{0}^{k})\Delta_{k,1}^{(2)}(s^{\prime})$ (under event $\cap_{k=\lceil K^{\gamma}\rceil}^{K-1}\mathcal{G}_{k}$)
where $\Delta_{k,h}^{(2)}(s)=\left(V_{h}^{k}(s)-V_{h}^{\bm{\pi}^{k}}(s)\right)\mathbbm{1}\{\bm{\pi}^{k}\notin\mathbf{\Pi}^{*}\}$, and the penultimate inequality follows from Hölder’s inequality, the definition of $Q_{1}^{k}$, and a similar argument as in (11). We now write the same set of inequalities for horizon $h$:
$\Delta_{k,h}^{(2)}(s_{h}^{k})$ $\leq 2b^{k}_{h}(n_{h}^{k}(s_{h}^{k},a_{h}^{k}))\mathbbm{1}\{\bm{\pi}^{k}\notin\mathbf{\Pi}^{*}\}+\mathbb{E}\left[\Delta_{k,h+1}^{(2)}(s_{h+1}^{k})\mid\mathcal{F}_{k,h}\right],$
where $\mathcal{F}_{k,h}=\sigma(\{s_{0}^{0},a_{0}^{0},s_{1}^{0},a_{1}^{0},\ldots,s_{H-1}^{0},a_{H-1}^{0},s_{0}^{1},a_{0}^{1},\ldots,s_{0}^{k},a_{0}^{k},\ldots,s_{h}^{k},a_{h}^{k}\})$ represents the history of states and actions observed by the algorithm up to episode $k$ and step $h$, and $a_{h}^{k}=\pi_{h}^{k}(s_{h}^{k})$. Next, we define a Martingale difference sequence $\{\eta_{k,h}^{(2)}(s_{h}^{k})\}$:
$\eta_{k,h}^{(2)}(s_{h}^{k})$ $=\Delta_{k,h}^{(2)}(s_{h}^{k})-\mathbb{E}\left[\Delta_{k,h}^{(2)}(s_{h}^{k})\mid\mathcal{F}_{k,h-1}\right].$
Rewriting $\Delta_{k,h}^{(2)}(s_{h}^{k})$ using the martingale difference sequence defined above, we have:
$\Delta_{k,h}^{(2)}(s_{h}^{k})-\Delta_{k,h+1}^{(2)}(s_{h+1}^{k})$ $\leq 2b^{k}_{h}(n_{h}^{k}(s_{h}^{k},a_{h}^{k}))\mathbbm{1}\{\bm{\pi}^{k}\notin\mathbf{\Pi}^{*}\}-\eta_{k,h+1}^{(2)}(s_{h+1}^{k}).$
Applying a telescoping sum over $h$ and noting that $\Delta_{k,H}^{(2)}\equiv 0$, we obtain the following upper bound:
$R_{K}^{(2)}$ $=\sum_{k=\lceil K^{\gamma}\rceil}^{K-1}\Delta_{k,0}^{(2)}(s_{0}^{k})\leq\sum_{k=\lceil K^{\gamma}\rceil}^{K-1}\sum_{h=0}^{H-1}2b^{k}_{h}(n_{h}^{k}(s_{h}^{k},a_{h}^{k}))\mathbbm{1}\{\bm{\pi}^{k}\notin\mathbf{\Pi}^{*}\}-\sum_{k=\lceil K^{\gamma}\rceil}^{K-1}\sum_{h=0}^{H-1}\eta_{k,h+1}^{(2)}(s_{h+1}^{k}).$
Notice that, under the event $\bigcap_{k=\lceil K^{\gamma}\rceil}^{K-1}\mathcal{G}_{k},$ the $\ell_{1}$-distance between the estimated and true transition kernels, $\hat{P}_{h}^{k}(\cdot\mid s,a)$ and $P_{h}^{*}(\cdot\mid s,a)$, becomes arbitrarily small whenever the visitation count $n_{h}^{k}(s,a)$ is sufficiently large. Moreover, when $n_{h}^{k}(s,a)$ is sufficiently large, the bonus terms become negligibly small. Consequently, the optimal policy sets for the MDPs with transition kernels and reward functions $(\bm{P}^{*},r)$ and $(\hat{\bm{P}}^{k},\,r+b^{k})$ coincide. More formally, by Lemma 2, under the event $\bigcap_{k=\lceil K^{\gamma}\rceil}^{K-1}\mathcal{G}_{k},$ for any episode $l$ such that
$n_{h}^{l}(s,a)>\bar{n}_{h}\coloneqq\frac{(2S\ln 2+4\mu K^{\alpha})(2H(H-h-1))^{2}}{(\texttt{gap}^{*})^{2}}\quad\text{for all }(s,a,h)\in\mathcal{S}\times\mathcal{A}\times[H],$
we have $\mathbbm{1}\{\bm{\pi}^{l}\notin\mathbf{\Pi}^{*}\}=0.$
We continue with bounding the summation $2\sum_{k=\lceil K^{\gamma}\rceil}^{K-1}\sum_{h=0}^{H-1}b^{k}_{h}(n_{h}^{k}(s_{h}^{k},a_{h}^{k}))\mathbbm{1}\{\bm{\pi}^{k}\notin\mathbf{\Pi}^{*}\}$. By definition, in either case of KI or KD, we have
$2\sum_{k=0}^{K-1}\sum_{h=0}^{H-1}b^{k}_{h}(n_{h}^{k}(s_{h}^{k},a_{h}^{k}))\mathbbm{1}\{\bm{\pi}^{k}\notin\mathbf{\Pi}^{*}\}$ (4)
$\qquad\leq 2\sum_{k=0}^{K-1}\sum_{h=0}^{H-1}(H-h-1)\mathbbm{1}\{\bm{\pi}^{k}\notin\mathbf{\Pi}^{*}\}\left(\sqrt{\frac{0.5S\ln 2+\mu K^{\alpha}}{\max\{1,n_{h}^{k}(s_{h}^{k},a_{h}^{k})\}}}\right)$
$\qquad\leq 2\sqrt{0.5S\ln 2+\mu K^{\alpha}}\sum_{k=0}^{K-1}\sum_{h=0}^{H-1}(H-h-1)\mathbbm{1}\{\bm{\pi}^{k}\notin\mathbf{\Pi}^{*}\}\sqrt{\frac{1}{\max\{1,n_{h}^{k}(s_{h}^{k},a_{h}^{k})\}}}$
$\qquad\leq 2\sqrt{0.5S\ln 2+\mu K^{\alpha}}\sum_{(s,a)\in\mathcal{S}\times\mathcal{A}}\sum_{h=0}^{H-1}(H-h-1)\sum_{n=1}^{\left\lfloor\bar{n}_{h}\right\rfloor}\sqrt{\frac{1}{n}}$
$\qquad\leq 4\sqrt{0.5S\ln 2+\mu K^{\alpha}}\sum_{(s,a)\in\mathcal{S}\times\mathcal{A}}\sum_{h=0}^{H-1}(H-h-1)\sqrt{\bar{n}_{h}}$
$\qquad\leq\frac{8HSA({S\ln 2+2\mu K^{\alpha}})}{\texttt{gap}^{*}}\sum_{h=0}^{H-1}(H-h-1)^{2}$
$\qquad\leq\frac{4H^{2}(H-1)(2H-1)SA({S\ln 2+2\mu K^{\alpha}})}{3\texttt{gap}^{*}}\eqqcolon m_{K}$ (5)
Since, under the good set, $V_{h}^{k}\geq V_{h}^{\bm{\pi}^{*}}\geq V_{h}^{\bm{\pi}^{k}}$, each $\Delta_{k,h}^{(2)}$ lies in $[0,H]$, which implies $|\eta_{k,h+1}^{(2)}|\leq H$. There are at most $HK$ such increments, so applying Azuma–Hoeffding yields
@@
$\mathbb{P}\left(\left\{R_{K}^{(2)}\geq x-\lceil K^{\gamma}\rceil H\right\}\cap(\cap_{k=\lceil K^{\gamma}\rceil}^{K-1}\mathcal{G}_{k})\right)\leq$ $\mathbb{P}\left(-\sum_{k=\lceil K^{\gamma}\rceil}^{K-1}\sum_{h=0}^{H-1}\eta_{k,h+1}^{(2)}(s_{h+1}^{k})\geq x-\lceil K^{\gamma}\rceil H-m_{K}\right)$

### v2

@@ 4 Main Results
It is natural to compare Corollary 1 with the existing gap-dependent bounds for tabular MDPs. For model-based algorithms, Simchowitz and Jamieson (2019) show that, given a global $Q^{\star}$-gap, for any fixed confidence level $\delta\in(0,1)$, the cumulative regret satisfies the following high-probability bound:
$\mathbb{P}\left(R_{K}\geq\tilde{\mathcal{O}}\!\Big(H^{4}SA(S\vee H)+\frac{H^{3}SA}{\mathrm{gap}^{*}}\log(K/\delta)\Big)\right)\;\leq\;\delta.$
@@
$\mathbb{E}[R_{K}]\;\leq\;\tilde{\mathcal{O}}\!\Big(\frac{H^{3}SA}{\mathrm{gap}^{*}}\log K\Big).$
Similar instance-dependent rates have been obtained for optimistic $Q$-learning and related model-free methods (Yang et al., 2021; Velegkas et al., 2022). All the algorithms presented in these works depend on the specific choice of confidence level $\delta\in(0,1)$.
In contrast, our results are instance-dependent in a comparable regime, but they are organized around the entire tail rather than a single quantile. Instead of working at a fixed confidence level $\delta$ and producing a bound, Corollary 1 describes how the entire curve $x\mapsto\mathbb{P}(R_{K}\geq x)$ behaves: for $x$ above the instance-dependent baseline $m_{K}^{\mathrm{mb}}$ we obtain a sub-Gaussian tail
$\mathbb{P}(R_{K}\geq x)\;\leq\;\exp\!\Big(-\tilde{\mathcal{O}}\big(x^{2}/(H^{5}K)\big)\Big),$
up to a transition scale that depends on $(H,K,\alpha)$, and beyond that scale the tail becomes sub-Weibull, with decay of order $\exp\{-\tilde{\mathcal{O}}(K^{\alpha})\}$ in the KD case or $\exp\{-\tilde{\mathcal{O}}(x^{\alpha}/H^{\alpha})\}$ in the KI case.
@@ 5.1 Proof of Theorem 1
Recall that $\bm{\pi}^{*}=\{\pi^{*}_{h}\}_{h=0}^{H-1}\in{\mathbf{\Pi}}^{*}$, where $\mathbf{\Pi}$$^{*}$ denotes the set of optimal policies for the transition kernel $\bm{P}^{*}=\{P^{*}_{h}\}_{h=0}^{H-2}$. At episode $k$, the algorithm computes a greedy policy $\pi^{k}=\{\pi^{k}_{h}\}_{h=0}^{H-1}$ using an optimistic estimate of the optimum value function $\{V^{*}_{h}\}_{h=0}^{H-1}$ based on data collected up to episode $k$. Recall that $R_{K}$ denote the cumulative regret incurred up to episode K1K-1. Then, writing $s_{0}^{k}\equiv s_{0}$, we have:
$R_{K}$ $=\sum_{k=0}^{K-1}\left(V_{0}^{*}(s_{0}^{k})-V_{0}^{\bm{\pi}^{k}}(s_{0}^{k})\right)=\underbrace{\sum_{k=0}^{\lceil K^{\gamma}\rceil-1}\left(V_{0}^{*}(s_{0}^{k})-V_{0}^{\bm{\pi}^{k}}(s_{0}^{k})\right)}_{R_{K}^{(0)}}+\underbrace{\sum_{k=\lceil K^{\gamma}\rceil}^{K-1}\left(V_{0}^{*}(s_{0}^{k})-V_{0}^{\bm{\pi}^{k}}(s_{0}^{k})\right)}_{R_{K}^{(1)}}.$
Here, $R_{K}^{(0)}\leq\lceil K^{\gamma}\rceil H$ denotes the initial regret incurred during episodes with insufficient data. We continue by analyzing $R_{K}^{(1)}$. Throughout this proof, we define the following clipping function:
$\mathrm{clip}(x)\;=\;\begin{cases}0,&\text{if }x<\mathrm{gap}^{*},\\[4.0pt] x,&\text{otherwise.}\end{cases}$
Note that for any non negative numbers $a_{1},a_{2},\dots,a_{H}$, the $\mathrm{clip}$ function satisfies
$\mathrm{clip}\left(\sum_{i=1}^{H}a_{i}\right)\leq\sum_{i=1}^{H}\mathrm{clip}(Ha_{i}).$ (4)
The regret associated with any episode $k\geq\left\lceil K^{\gamma}\right\rceil$ is given by
$V_{0}^{*}(s_{0}^{k})-V_{0}^{\bm{\pi}^{k}}(s_{0}^{k})$ $=V_{0}^{*}(s_{0}^{k})-Q_{0}^{*}(s_{0}^{k},a_{0}^{k})+(Q_{0}^{*}(s_{0}^{k},a_{0}^{k})-Q_{0}^{\bm{\pi}^{k}}(s_{0}^{k},a_{0}^{k}))$
$=\mathrm{gap}_{0}(s_{0}^{k},a_{0}^{k})+\mathbb{E}\left[V_{1}^{*}(s_{1}^{k})-V_{1}^{\bm{\pi}^{k}}(s_{1}^{k})\big|\mathcal{F}_{0}^{k}\right]$
where $\mathcal{F}_{h}^{k}=\sigma(\{s_{0}^{0},a_{0}^{0},s_{1}^{0},a_{1}^{0},\ldots,s_{H-1}^{0},a_{H-1}^{0},s_{0}^{1},a_{0}^{1},\ldots,s_{0}^{k},a_{0}^{k},\ldots,s_{h}^{k},a_{h}^{k}\})$ represents the history of states and actions observed by the algorithm up to episode $k$ and step $h$, and $a_{h}^{k}=\pi_{h}^{k}(s_{h}^{k})$. Define the value difference $\Delta_{h}^{k}(s)\coloneqq V_{h}^{*}(s)-V_{h}^{\bm{\pi}^{k}}(s)$. Following a similar argument as above, for all $h\in[H]$, we have
$\Delta^{k}_{h}(s_{h}^{k})$ $=\mathrm{gap}_{h}(s_{h}^{k},a_{h}^{k})+\mathbb{E}\left[\Delta^{k}_{h+1}(s_{h+1}^{k})\big|\mathcal{F}_{h}^{k}\right]$
$=\mathrm{gap}_{h}(s_{h}^{k},a_{h}^{k})+\Delta^{k}_{h+1}(s_{h+1}^{k})-\eta^{k}_{h+1}(s_{h+1}^{k}),$
where the sequence $\{\eta^{k}_{h}(s_{h}^{k})\}_{h=1,k=0}^{H-1,K-1}$ defined as
$\eta^{k}_{h}(s_{h}^{k})=\Delta^{k}_{h}(s_{h}^{k})-\mathbb{E}\left[\Delta^{k}_{h}(s_{h}^{k})\big|\mathcal{F}_{h-1}^{k}\right]$
is a Martingale difference sequence with the filtration $\{\mathcal{F}_{h-1}^{k}\}_{h=1,k=0}^{H-1,K-1}$. For the sake of notational simplicity, we define $\eta^{k}_{H}\equiv 0$ for all $k\in[K]$. Using a telescopic sum, noting that $\Delta^{k}_{H}\equiv 0$, we obtain
$\Delta^{k}_{0}(s_{0}^{k})$ $=\sum_{h=0}^{H-1}\mathrm{gap}_{h}(s_{h}^{k},a_{h}^{k})-\sum_{h=0}^{H-1}\eta^{k}_{h+1}(s_{h+1}^{k})$
$=\sum_{h=0}^{H-1}\mathrm{clip}\left(\mathrm{gap}_{h}(s_{h}^{k},a_{h}^{k})\right)-\sum_{h=0}^{H-1}\eta^{k}_{h+1}(s_{h+1}^{k})$ (5)
where the last equality follows by the fact that $\mathrm{gap}_{h}(s,a)\geq\mathrm{gap}^{*}$ if $a\notin\mathcal{A}^{*}_{h}(s)$ and $\mathrm{gap}_{h}(s,a)=0$ otherwise.
@@
Note that $\mathrm{gap}_{h}(s_{h}^{k},a_{h}^{k})=V_{h}^{*}(s_{h}^{k})-Q_{h}^{*}(s_{h}^{k},a_{h}^{k})\leq V_{h}^{*}(s_{h}^{k})-V_{h}^{k}(s_{h}^{k})+V_{h}^{k}(s_{h}^{k})-V_{h}^{\bm{\pi}^{k}}(s_{h}^{k})$. The bonus function $b_{h}^{k}(\cdot)$ is chosen so that $V_{h}^{k}$ provides an optimistic estimate of $V_{h}^{*}$, as long as the estimated transition kernel $\hat{P}_{h^{\prime}}^{k}$ and the true transition kernel $P_{h^{\prime}}^{*}$ remain close for all horizons $h^{\prime}\geq h$. This combined with the monotonicity of the $\mathrm{clip}(\cdot)$ function provides an upper bound for the regret at episode $k\geq\lceil K^{\gamma}\rceil$. We define the “good event at episode kk” as
$\mathcal{G}^{k}=$ $\bigcap_{h=0}^{H-2}\bigcap_{s,a}\left\{\|P_{h}^{*}(\cdot|s,a)-\hat{P}_{h}^{k}(\cdot|s,a)\|_{1}\leq\frac{2b_{h}^{k}(n_{h}^{k}(s,a))}{H-h-1}\right\}.$
@@
Under the event $\mathcal{G}^{k}$, we claim that $Q_{h}^{k}(s,a)\geq Q_{h}^{*}(s,a)$ for all $(s,a,h)\in\mathcal{S}\times\mathcal{A}\times[H]$. We establish this by induction on the horizon. At the terminal horizon, we have $Q_{H-1}^{k}(s,a)=Q_{H-1}^{*}(s,a)$ for all $(s,a)\in\mathcal{S}\times\mathcal{A}$. Assuming $Q_{h+1}^{k}(s,a)\geq Q_{h+1}^{*}(s,a)$ for all $(s,a)$, we now show that the same inequality holds for horizon $h$.
@@
$\qquad\geq-b_{h}^{k}(n_{h}^{k}(s,a)),$ (by definition of ${\mathcal{G}}^{k}$)
where $c=\tfrac{1}{2}\big(\max_{x}V_{h+1}^{*}(x)+\min_{x}V_{h+1}^{*}(x)\big),$ $\mathrm{span}(f)=\max_{x}f(x)-\min_{x}f(x)$, and $\mathbf{1}$ denotes the all-ones vector. Hence, under $\bar{\mathcal{G}}^{K}$, taking maxima and recursing yields
$Q_{h}^{k}(s,a)\geq Q_{h}^{*}(s,a),\qquad V_{h}^{k}(s)\geq V_{h}^{*}(s),$ (6)
for all $(s,a,h)\in\mathcal{S}\times\mathcal{A}\times[H]$.
Defining $\bar{\mathcal{G}}^{K}\coloneqq\cap_{k=\lceil K^{\gamma}\rceil}^{K-1}\mathcal{G}^{k}$, we next bound the probability of the bad event $\left(\bar{\mathcal{G}}^{K}\right)^{\mathrm{c}}$:
$\mathbb{P}\left(\left(\bar{\mathcal{G}}^{K}\right)^{\mathrm{c}}\right)$
@@
$\leq\sum_{s,a,h}\sum_{n=1}^{K}2^{S}\exp\left(-\frac{n}{2}\left(\frac{2b_{h}^{n\vee\lceil K^{\gamma}\rceil}(n)}{H-h-1}\right)^{2}\right)$ ((Weissman et al., 2003, Theorem 2.1))
$\leq SAHK\begin{cases}\displaystyle\ \exp\left(-2\mu K^{\alpha}\right),&\text{{KD}},\\[8.61108pt] \displaystyle\ \exp\left(-2\mu K^{\gamma\alpha}\right),&\text{{KI}}.\end{cases}\eqqcolon\delta_{K}^{\mathrm{mb}}(\gamma)$ (Lemma 1)
@@
Decomposing the tail probability of the regret into the contribution of the bad event and its complement, we have
$\mathbb{P}(R_{K}\geq x)$ $\leq\mathbb{P}(R_{K}^{(1)}\geq x-\lceil K^{\gamma}\rceil H)$
@@
$\leq\mathbb{P}\left(\left(\bar{\mathcal{G}}^{K}\right)^{\mathrm{c}}\right)+\mathbb{P}\left(\left\{R_{K}^{(1)}\geq x-\lceil K^{\gamma}\rceil H\right\}\cap\bar{\mathcal{G}}^{K}\right)$
@@
$\leq\delta_{K}^{\mathrm{mb}}(\gamma)+\mathbb{P}\left(\left\{R_{K}^{(1)}\geq x-\lceil K^{\gamma}\rceil H\right\}\cap\bar{\mathcal{G}}^{K}\right)$